# LinguaLIFT: An Effective Two-stage Instruction Tuning Framework for Low-Resource Language Tasks

### v1

@@ A.4 Training Details
We use LlamaFactory [^7] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024) as the training codebase for our experiments. In the first stage, we train only the additional alignment modules (e.g., the language alignment layer and the boundary tokens) for 3 epochs, with a constant learning rate of 6e-4 and a batch size of 256. In the second stage, we fine-tune all parameters of the LLM for 3 epochs. The learning rate is set to 2e-5, with a warm-up ratio of 0.05 and a cosine learning rate scheduler. We also apply a weight decay of 1e-2 and use a batch size of 128. All experiments are conducted on eight NVIDIA A100 GPUs for a day.
@@ Appendix B Details of Constructing the MMWP Benchmark
B.1 Selection of Target Languages and Categorization of Resource-level.
@@
B.2 Quality Estimation proves the effectiveness of the proposed MMWP benchmark.
To ensure the quality of the proposed MMWP benchmark, we conducted both human and automatic quality estimation evaluations.
For human evaluation, we employed five annotators, each of whom was responsible for post-calibrating and assessing the quality of the translated dataset. To facilitate this, we utilized GPT-4 for back-translation of the problems into English, a process that allowed the annotators to compare the translated versions with their original counterparts. The evaluators were tasked to scrutinize the translations for grammatical correctness, fluency, and semantic accuracy. Any significant deviations in meaning between the original and translated versions were carefully addressed. The annotators were guided to revise such instances to better align with the original intent, thereby ensuring the preservation of the original linguistic integrity. Moreover, the post-editing process was carried out to maintain native language properties, such as syntactical flexibility and natural phrasing, thus ensuring the fluency and native-like quality of the back-translations. This rigorous human calibration process aimed to certify that the final MMWP dataset was linguistically accurate and culturally appropriate for all represented languages.
The prompt for back-translating the low-resource language texts and assisting the post-editing process are presented below.
@@ 3. Semantic Accuracy:
[…]
Average 78.71 86.08 1.09
The results from both the human and automatic evaluations indicate that the dataset is properly constructed and adequately reflects the multilingual nature of the tasks. The human post-editing process ensured that the translated problems maintained their semantic integrity, while the automatic evaluation confirmed that the translations preserved the necessary linguistic structure for multilingual reasoning. Overall, the quality assurance measures we implemented guarantee that the MMWP benchmark is reliable and effective in evaluating multilingual mathematical reasoning across diverse languages and resource levels.
@@ C.1 Evaluation Results on MMWP
The complete experimental results on MMWP are shown in Table 7, 8 and 9. These tables present Mono-SFT, Multi-SFT, and Leveraging External Tools or Models category comparison baselines.
@@ C.2 Evaluation Results on MGSM and MMWP
The complete experimental results on MGSM and MSVAMP are shown in Table 10, and 11.
@@
LinguaLIFT-MetaMath 54.4 62.0 64.8 63.6 56.8 50.0 60.4 55.6 54.0 57.6 54.0 55.4 58.8 57.6
LinguaLIFT-OpenMath2 63.0 69.6 76.0 75.2 62.8 52.0 71.6 64.2 61.2 66.6 58.0 63.8 66.5 65.5
Table 10: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Yoon et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The dagger symbol (†) indicates results obtained using officially released models, while the double dagger symbol (‡) denotes results from our local implementation.
@@
LinguaLIFT-MetaMath 54.3 62.8 65.2 59.9 62.7 57.9 59.7 56.8 57.2 56.0 56.1 60.6 59.3
LinguaLIFT-OpenMath2 65.2 76.4 77.9 74.2 74.5 73.5 71.4 68.6 67.9 72.2 67.2 74.3 72.2
@@ C.3 Evaluation Results on XNLI and X-CSQA
The complete experimental results on XNLI and X-CSQA are shown in Table 10, and 11.
@@ D.1 Ablation of Two-Stage Training
To demonstrate the necessity of the proposed two-stage instruction tuning approach, we perform an ablation study on the distinct components of the two-stage process: the Language-Align stage and the Task-Transfer stage. This study is conducted on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of each stage’s removal on reasoning performance across all languages. The detailed results are presented in Tables 14 and 15 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of the specific stage.
@@
When the Task-Transfer stage was removed, we fine-tuned the language alignment layer using only code-switched translation data without updating the LLM’s parameters. After ablating the Task-Transfer stage, the models couldn’t complete reasoning tasks across all languages, resulting in an average accuracy of just 8.49% on the MGSM dataset and 9.74% on the MSVAMP dataset. This demonstrates that task-specific instruction tuning is critical for enabling the LLM to perform task-solving operations effectively.
@@ D.2 Ablation of Trainable Modules
To further demonstrate the essential design of the two-stage training approach, we conduct an ablation study on the trainable modules during each stage. This study is performed on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of training each module at the appropriate stage. The detailed results are presented in Tables 16 and 17 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of a specific operation.
@@
D.3 Analysis of Code-Switch Tuning
Part-of-Speech #Nums
Noun 84726
Verb 42451
Adposition 37727
Pronoun 25012
Adjective 21790
Conjunction 20764
Auxiliary 16409
Adverb 9150
Appendix E Supplementary Experiments
(a) MGSM
(b) MSVAMP
Test set MGSM MSVAMP
$\rho$ 0.733* 0.778* 0.763* 0.783*
E.2 Language Transferability in Language Families and Writing Systems
Figure 8: The experimental results grouped by diverse language families and writing systems of different models.
@@ E.2 Language Transferability in Language Families and Writing Systems
E.3 Adapting to different types and scales of LLMs
LinguaLIFT can be flexibly adapted to different LLMs. To validate this, we conducted experiments on Mistral-7B (Jiang et al., 2023) and the larger Llama-2-13B (Touvron et al., 2023). As shown in Tables 22 and 23, LinguaLIFT outperforms various baselines, with average accuracy improvements of at least 1.6% and 0.9% for low-resource languages based on Llama-2-13B and Mistral-7B, respectively. These results demonstrate LinguaLIFT’s potential for broader applicability across LLMs.
Llama-2-13B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
Multi-SFT
Leveraging External Tools or Models
our methods
Table 22: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (citation: LLaMA2), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024b); Huang et al. (2024).
Mistral-7B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
Multi-SFT
Leveraging External Tools or Models
our methods
Table 23: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024b); Huang et al. (2024).
E.4 Impact of Different Language Alignment Layers
@@
E.5 Impact of Multilingual Encoder Size on Language Alignment
Figure 9: MGSM accuracy(%) of LinguaLIFT models using different kinds and types of multilingual encoders.
E.6 Incorporating multilingual reasoning data in LinguaLIFT
Figure 10: Effects of tuning LLM with mixed supervised data. Generally, incorporating multilingual supervised data into LinguaLIFT can achieve a higher ceiling for low-resource language tasks performance.
E.7 LinguaLIFT CoT Examples
In this section, we present three examples of CoT reasoning generated by LinguaLIFT models. These examples demonstrate the ability of LinguaLIFT to effectively understand low-resource languages and generate corresponding mathematical reasoning processes using the chain-of-thought reasoning strategy.
Figure 11 illustrates a zero-shot CoT example from the MGSM Bengali dataset. In this case, the vanilla Mono-SFT model produces incorrect reasoning, while the LinguaLIFT model generates a correct reasoning process, ultimately leading to the correct answer. Similar trends are observed in the Thai and Swahili examples, shown in Figures 12 and 13, where the LinguaLIFT model corrects the erroneous reasoning produced by the Mono-SFT model when presented with the same mathematical problems.

### v2

@@ A.4 Training Details
We use mT5 (Xue et al., 2021) as the pre-trained multilingual encoder illustrated in Section §3. We extend LlamaFactory [^4] https://github.com/hiyouga/LLaMA-Factory (Zheng et al., 2024) as the training codebase for our experiments. In the first stage, we train only the additional alignment modules (e.g., the language alignment layer and the boundary tokens) for 3 epochs, with a constant learning rate of 6e-4 and a batch size of 256. In the second stage, we fine-tune all parameters of the LLM for 3 epochs. The learning rate is set to 2e-5, with a warm-up ratio of 0.05 and a cosine learning rate scheduler. We also apply a weight decay of 1e-2 and use a batch size of 128. All experiments are conducted on eight NVIDIA A100 GPUs for a day.
@@ Appendix B Details of Constructing the MMWP Benchmark
B.1 Language Categorization List.
  • Low-resource languages: Afrikaans (af), Arabic (ar), Belarusian (be), Bengali (bn), Basque (eu), Gujarati (gu), Hausa (ha), Hindi (hi), Armenian (hy), Icelandic (is), Kannada (kn), Luxembourgish (lb), Macedonian (mk), Malayalam (ml), Marathi (mr), Nepali (ne), Slovak (sk), Swahili (sw), Tamil (ta), Telugu (te), Thai (th).
  • Medium-resource languages: Bulgarian (bg), Catalan (ca), Czech (cs), Danish (da), Finnish (fi), Croatian (hr), Hungarian (hu), Indonesian (id), Korean (ko), Norwegian Bokmål (nb), Polish (pl), Portuguese (pt), Romanian (ro), Slovenian (sl), Serbian (sr), Ukrainian (uk), Vietnamese (vi).
  • High-resource languages: German (de), English (en), Spanish (es), French (fr), Italian (it), Japanese (ja), Dutch (nl), Russian (ru), Swedish (sv), Chinese (zh).
B.2 Selection of Target Languages and Categorization of Resource-level.
@@
B.3 Evaluation and post-calibration process proves the reliability of the MMWP benchmark.
To ensure the quality of the proposed MMWP benchmark, we conducted both human and automatic quality estimation evaluations and further post-calibration processes.
For human evaluation, we employed a total of five annotators for all languages, each responsible for assessing the quality of the translated dataset and performing post-editing and calibration on the translations to ensure consistency and quality across the dataset. These annotators, native Chinese speakers with professional English proficiency and extensive translation and linguistic research experience, reviewed translations for grammatical correctness, fluency, and semantic accuracy. To ensure the reliability of the annotations, we hired two experts who are from Lan-bridge, a qualified institution for translation services [^5] Requirements for translation services: https://www.iso.org/standard/59149.html., recognized by the ISO [^6] International Organization for Standardization: https://www.iso.org/home.html., to serve as instructors to assess the reliability of the annotation. They evaluated the randomly selected translations and reannotated those with significant deviations. The evaluation process also included back-translation using GPT-4, which enabled annotators to compare the translated versions with the original texts. Any substantial deviations were addressed through post-editing to align the translations with the original intent and maintain native-like fluency. We carefully measured inter-annotator agreement (IAA) to ensure consistency and reliability. We found substantial agreement among annotators, with a Cohen Kappa score of 0.67 on randomly selected 50 instances, indicating substantial consistency in evaluating the translations. In cases where errors were identified, the annotators worked together to correct them, with an improved IAA of 0.80. The two experts’ determinations are consistent, with an IAA of 0.78.
[…]
Average 78.71 86.08 1.09
The results from both the human and automatic evaluations indicate that the dataset is properly constructed and adequately reflects the multilingual nature of the tasks. The human post-editing process ensured that the translated problems maintained their semantic integrity, while the automatic evaluation confirmed that the translations preserved the necessary linguistic structure for multilingual reasoning. Overall, the quality assurance measures we implemented to guarantee that the MMWP benchmark is reliable and effective in evaluating multilingual mathematical reasoning across diverse languages and resource levels.
The prompts for assisting the post-editing process and back-translating the low-resource language texts are presented below.
@@ C.1 Evaluation Results on MMWP
The complete experimental results on MMWP are shown in Table 8, 9 and 10. These tables present Mono-SFT, Multi-SFT, and Leveraging External Tools or Models category comparison baselines.
@@ C.2 Evaluation Results on MGSM and MMWP
The complete experimental results on MGSM and MSVAMP are shown in Table 11, and 12.
@@
LinguaLIFT-MetaMathQA 54.4 62.0 64.8 63.6 56.8 50.0 60.4 55.6 54.0 57.6 54.0 55.4 58.8 57.6
[…]
LinguaLIFT-OpenMathInstruct-2 63.0 69.6 76.0 75.2 62.8 52.0 71.6 64.2 61.2 66.6 58.0 63.8 66.5 65.5
LinguaLIFT-OpenMathInstruct-2$\Diamond$ 59.2 65.0 72.8 71.6 59.4 48.6 67.2 60.0 57.4 62.8 55.4 59.9 62.9 61.8
Table 11: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Yoon et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The dagger symbol (†) indicates results obtained using officially released models, the diamond symbol ($\Diamond$) signifies the LoRA implementation, and the double dagger symbol (‡) denotes results from our local implementation.
@@
LinguaLIFT-MetaMathQA 54.3 62.8 65.2 59.9 62.7 57.9 59.7 56.8 57.2 56.0 56.1 60.6 59.3
[…]
LinguaLIFT-OpenMathInstruct-2 65.2 76.4 77.9 74.2 74.5 73.5 71.4 68.6 67.9 72.2 67.2 74.3 72.2
LinguaLIFT-OpenMathInstruct-2$\Diamond$ 61.3 72.1 74.2 70.5 70.8 69.7 67.5 64.9 64.2 68.5 63.5 70/5 68.4
@@ C.3 Evaluation Results on XNLI and X-CSQA
The complete experimental results on XNLI and X-CSQA are shown in Table 13, and 14.
@@ D.1 Ablation of Two-Stage Training
To demonstrate the necessity of the proposed two-stage instruction tuning approach, we perform an ablation study on the distinct components of the two-stage process: the Language-Align stage and the Task-Transfer stage. This study is conducted on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of each stage’s removal on reasoning performance across all languages. The detailed results are presented in Tables 15 and 16 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of the specific stage.
@@
When the Task-Transfer stage was removed, we fine-tuned the language alignment layer using only code-switched translation data without updating the LLM’s parameters. After ablating the Task-Transfer stage, the models couldn’t complete reasoning tasks across all languages, resulting in an average accuracy of just 8.49% on the MGSM dataset and 9.74% on the MSVAMP dataset. This demonstrates that reasoning instruction tuning is critical for enabling the LLM to perform reasoning operations effectively.
@@ D.2 Ablation of Trainable Modules
To further demonstrate the essential design of the two-stage training approach, we conduct an ablation study on the trainable modules during each stage. This study is performed on the mathematical reasoning datasets MGSM and MSVAMP to assess the impact of training each module at the appropriate stage. The detailed results are presented in Tables 17 and 18 for MGSM and MSVAMP, respectively, where “w/o” indicates the absence of a specific operation.
@@
D.3 Impact of Multilingual Encoder Size on Language Alignment
D.4 Adapting to different types and scales of LLMs
LinguaLIFT can be flexibly adapted to different LLMs. To validate this, we conducted experiments on Mistral-7B (Jiang et al., 2023) and the larger Llama-2-13B (Touvron et al., 2023). As shown in Tables 19 and 20, LinguaLIFT outperforms various baselines, with average accuracy improvements of at least 1.6% and 0.9% for low-resource languages based on Llama-2-13B and Mistral-7B, respectively. These results demonstrate LinguaLIFT’s potential for broader applicability across LLMs.
Llama-2-13B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
Multi-SFT
Leveraging External Tools or Models
our methods
Table 19: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (citation: LLaMA2), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024); Huang et al. (2024).
Mistral-7B as base model Bn De En Es Fr Ja Ru Sw Th Zh LR. HR. Avg.
baseline
Mono-SFT
Multi-SFT
Leveraging External Tools or Models
our methods
Table 20: Experimental Results on the MGSM Dataset. "LR." "HR." and "Avg." represent the average performance for low-resource languages, high-resource languages, and all languages, respectively. Following prior work (Huang et al., 2024), we classify Bn, Te, Th, and Sw as low-resource languages, while the remaining languages are categorized as high-resource. The asterisk symbol (*) denotes results taken directly from the published results of Zhu et al. (2024); Huang et al. (2024).
D.5 Language Transferability in Language Families and Writing Systems
@@
D.6 Analysis of Code-Switch Tuning
Part-of-Speech #Nums
Noun 84726
Verb 42451
Adposition 37727
Pronoun 25012
Adjective 21790
Conjunction 20764
Auxiliary 16409
Adverb 9150
Figure 11: Effects of tuning LLM with mixed supervised data. Generally, incorporating multilingual supervised data into LinguaLIFT can achieve a higher ceiling for low-resource language reasoning performance.
D.7 Incorporating multilingual reasoning data in LinguaLIFT
Appendix E Supplementary Experiments
(a) MGSM
(b) MSVAMP
@@ E.1 Quantitative Analysis of the Correlation between Multilingual Alignment and Reasoning Performance
Test set MGSM MSVAMP
$\rho$ 0.733* 0.778* 0.763* 0.783*
E.2 Impact of Different Language Alignment Layers
E.3 LinguaLIFT CoT Examples
We present three examples of CoT reasoning. These examples demonstrate LinguaLIFT’s ability to effectively understand low-resource languages and generate corresponding mathematical reasoning processes using the chain-of-thought reasoning strategy. Figure 13 illustrates a zero-shot CoT example from the MGSM Bengali dataset. The Vanilla Mono-SFT model produces incorrect reasoning, while the LinguaLIFT model generates a correct reasoning process, ultimately leading to the correct answer. Similar trends are observed in the Thai and Swahili, shown in Figures 14 and 15, where the LinguaLIFT model corrects the erroneous reasoning produced by the Mono-SFT model when presented with the same problems.